OpenStudy (anonymous):

3v^2-20x-32 _______________ 2x^2-x-15 __________________ 9x^2+24x+16 ________________ 2x^2+7x+5

OpenStudy (anonymous):

does anybody out there understand this?

OpenStudy (mathteacher1729):

Is that all one gigantic fraction? \[ \huge \dfrac{\left( \frac{3v^2-20x-32}{2x^2-x-15}\right) }{\left( \frac{9x^2+24x+16}{2x^2+7x+5}\right) } \]

OpenStudy (anonymous):

yes

OpenStudy (mathteacher1729):

Goodness gracious, you got a lot of factoring to do. :-p FIRST though, "invert & Multiply. \[ \huge \left( \frac{3v^2-20x-32}{2x^2-x-15}\right) \cdot \left( \frac{2x^2+7x+5}{9x^2+24x+16} \right) \]

OpenStudy (mathteacher1729):

NOW factor each expression individually. Can you factor \[ \huge 3v^2-20x -32\] for instance?

OpenStudy (anonymous):

I am a bit confused by it all

OpenStudy (anonymous):

yes, and thats supposed to be an x

OpenStudy (mathteacher1729):

This can certainly be overwhelming if you're not familiar with four main things: 1. Simplifying simple fractions. (with numbers, no variables) 2. Simplifying simple fractions (with variables) 3. Factoring quadratic equations 4. Simplifying complex fractions (with numbers, no variables)

OpenStudy (anonymous):

yes i am in over my head but trying to figur it out

OpenStudy (anonymous):

ok so if I get it all factored, whats next?

OpenStudy (mathteacher1729):

You go cancel-crazy. If you have the same term in the numerator and denominator, cancel it. But first, let's talk about factoring quadratic equations... 3v^2-20x-32 Can you factor that? Just, focus on that only right now. :)

OpenStudy (anonymous):

(3x+8 )(x-4 )??

OpenStudy (mathteacher1729):

The x terms don't work out. 8x -12x = -4x

OpenStudy (anonymous):

oh (3x+4)(x-8)

OpenStudy (mathteacher1729):

Yes, you got it. :) 3x*x = 3x^2 4x -24x = -20x 4*(-8) = -32

OpenStudy (mathteacher1729):

Ok, now, can you factor 2x^2-x-15

OpenStudy (anonymous):

so is the key to this whole factoring thing just trial and error? or is there some magic concept that can be used?

OpenStudy (anonymous):

(2x+ )(x- ) I honestly dont know what I should use here

OpenStudy (mathteacher1729):

It's just practice, really. You start to spot patterns. That's a good start. Next step = What are the factors of 15 ?

OpenStudy (anonymous):

5&3 or 15&1

OpenStudy (mathteacher1729):

Yup. So the key is experimenting with those. The answer is somewhere in there. Just takes some playing around with. :)

OpenStudy (anonymous):

(2x-1)(x+15)?

OpenStudy (mathteacher1729):

The x terms don't work out -1x + 30x = 29x

OpenStudy (anonymous):

yeah this is where i get confused a bit... the -1 makes the last part right but not the middle number

OpenStudy (mathteacher1729):

Hint: 1 and 15 won't do it. No matter how you try, it won't work.

OpenStudy (anonymous):

ok...so (2x+3)(x-5)

OpenStudy (anonymous):

ahh (2x+5)(x-3)

OpenStudy (anonymous):

hey thanks for your help and your patience! i just dont get this stuff... my brain doesnt work this way... i will probably never understand it

OpenStudy (mathteacher1729):

Bingo.

OpenStudy (anonymous):

ok so the 2nd part i just do the same thing... but then what? do i have to invert something there?

OpenStudy (mathteacher1729):

So the key is to factor all four quadratics. I already "flipped and multiplied". \[\huge \left( \frac{3v^2-20x-32}{2x^2-x-15}\right) \cdot \left( \frac{2x^2+7x+5}{9x^2+24x+16} \right)\] Then when you have everything factored, cancel whatever you can. \[\huge \frac{( )( )}{()()}\cdot \frac{()()}{()()}\]

OpenStudy (anonymous):

ok thanks a bunch!

OpenStudy (mathteacher1729):

You're welcome! :) Have a good evening.

OpenStudy (anonymous):

you too :)